OpenStudy (tester97):

what is the difference between the commutative property of multiplication and the associative property of multiplication?

OpenStudy (anonymous):

oh I have this in my notes somewhere

sam (.sam.):

Commutative property: When two numbers are multiplied together, the product is the same regardless of the order of the multiplicands. For example 4 * 2 = 2 * 4 Associative Property: When three or more numbers are multiplied, the product is the same regardless of the grouping of the factors. For example (2 * 3) * 4 = 2 * (3 * 4) Associative indicates "group numbers" multiplying and changing of position of numbers

OpenStudy (anonymous):

GG, sam answered it like a boss.

OpenStudy (tester97):

true dat